OpenStudy (anonymous):

Two coils having equal resistance but different inductances are connected in series, then the time constant of series combination will ______ a. sum of time constants of individual coils b. average of time constant of individual coils c. geometric mean of the time constants of individual coils d. product of the time constants of individual coils... please help me with the answer and tell me how...

OpenStudy (anonymous):

could anyone help me with this question please..

OpenStudy (anonymous):

It's answer should be (b) average of time constant of individual coil.As time constant of a circuit would be L1+L2/R1+R2 but R1=R2=R so the equation become (L1+L2)/2R which is the required time constant and option (b) satisfies it as it says average of two time constant which is T=(T1+T2)/2 T1=L1/R1 =L1/R And, T2=L2/R2 =L2/R so, Avg =( L1/R +L2/R)/2 =(L1+L2)/2R I hope it will work for you.